A single phase transformer has 480 turns on the primary winding and 90 turns on the secondary winding. The maximum value of the magnetic flux density is 1.1T when 2200 volts, 50Hz is applied to the transformer primary winding. Calculate maximum flux, cross sectional area of core and secondary induced emf


1
Expert's answer
2021-05-19T06:14:36-0400

a. E="4.44* \phi fN"

"2200=4.44* \phi *480 *50"

"\phi=\frac{2200}{4.44*480*50}=0.0206456Weber"

b. "d=BA \implies A=\frac{d}{B}=\frac{0.0206456}{1.1}=0.0187687 m^2"

c. "\frac{E_2}{E_1}=\frac{N_2}{N_1} \implies E_2=\frac{N_2}{N_1} *E_1\implies E_2=\frac{90}{480} *2200=412.5 V"